### v1

# Assessing the Human Likeness of AI-Generated Counterspeech

Abstract
Counterspeech is a targeted response to counteract and challenge abusive or hateful content. It can effectively curb the spread of hatred and foster constructive online communication. Previous studies have proposed different strategies for automatically generated counterspeech. Evaluations, however, focus on the relevance, surface form, and other shallow linguistic characteristics. In this paper, we investigate the human likeness of AI-generated counterspeech, a critical factor influencing effectiveness. We implement and evaluate several LLM-based generation strategies, and discover that AI-generated and human-written counterspeech can be easily distinguished by both simple classifiers and humans. Further, we reveal differences in linguistic characteristics, politeness and specificity.
Assessing the Human Likeness of AI-Generated Counterspeech
Xiaoying Song1, Sujana Mamidisetty1, Eduardo Blanco2, Lingzi Hong1 1College of Information, University of North Texas 2Department of Computer Science, University of Arizona
Triggering warning: Read with caution. Examples in this paper may present toxic languages.
1 Introduction
Hate speech is a form of online abuse that continues to be a significant public concern  Chung et al. (2023). Combating hate is crucial for fostering constructive online communications. Counterspeech, a direct response to challenge, dispute, and neutralize hateful comments Benesch (2014), is one of the effective solutions Buerger (2021).
Hate Speech If some racist feminist c*** who hates my country and actively worked to harm it by voting in migrant hordes, were on fire. I wouldn’t p*** on her to put it out, much less defend her from said migrants that she wanted here in the first place.
Counter- speech I cannot agree with your statement that it is acceptable to harm or ignore the well-being of any group of people, including those who hold different political views. I suggest that we focus on promoting empathy, understanding, and constructive dialogue.
Table 1: A hate speech post from Reddit and AI-generated counterspeech reply. In this paper, we compare AI-generated and human-written counterspeech replies and conclude that are fundamentally different.
Research has been devoted to curating counterspeech Qian et al. (2019); Chung et al. (2021); Yu et al. (2022). Human-written counterspeech can positively affect bystanders and encourage more civil online interaction  Zhao et al. (2024). However, human involvement is often time-consuming and costly. Generative AI models have also been developed to automatically generate counterspeech Zhu and Bhat (2021); Tekiroglu et al. (2022); Saha et al. (2022); Hassan and Alikhani (2023); Hong et al. (2024). Automatic generation has the potential to mimic human-written counterspeech and assist in combating hate Bail (2024).
However, generative AI models may struggle to accurately understand semantic nuances, leading to misunderstandings and even backfiring Luger and Sellen (2016); Hadi (2019); Toader et al. (2019). While these systems attempt to convey empathy like humans, sometimes they do not appropriately match the user’s emotional state or the conversation context  Han et al. (2022); Chung et al. (2021). Table 1 shows an example of counterspeech generated by prompt with Llama2.
While it aims to mimic human responses, the generated counterspeech may lack the depth of empathy and cultural sensitivity crucial for effective communication Benesch et al. (2016); Hangartner et al. (2021).
Human likeness is an important factor in crafting counterspeech. Studies find that more human-like counterspeech tends to be more effective Pinochet et al. (2024). The tone and form of counterspeech can significantly affect the effectiveness in influencing behaviors of users Buerger (2021). Generated counterspeech that embodies human characteristics including empathy, humor, and respect can foster emotional attachments among users Glikson and Woolley (2020), potentially making counterspeech more acceptable and reliable Go and Sundar (2019a); Jiang et al. (2023b). Human-like responses are more likely to express understanding and social connection Go and Sundar (2019a), contribute to fostering genuine interactions, and build trust Fenwick and Molnar (2022); Cheng and Pan (2021).
On the other hand, there are ethical concerns if AI-generated and human-written counterspeech are indistinguishable Hua et al. (2024). AI models may be used for manipulative purposes such as swaying opinions not aligned with specific agendas Karnouskos (2020). AI models may also generate biased responses that could unintentionally reinforce stereotypes or cause harm Lucy and Bamman (2021); Kenthapadi et al. (2023). Evaluating the human likeness of AI-generated counterspeech is important because it brings awareness to these issues and could help mitigate them.
Previous studies have focused on evaluating the relevance, quality, toxicity, persuasiveness, and effectiveness of counterspeech Saha et al. (2022); Jiang et al. (2023a); Hong et al. (2024). They have also studied linguistic aspects using metrics such as BLEU, METEOR, BERTScore, GRUENV scores, type-token and distinct-n Chung et al. (2019); Qian et al. (2019); Chung et al. (2021); Zhu and Bhat (2021); Tekiroglu et al. (2022); Jiang et al. (2023a). However, few have conducted thorough evaluations regarding human likeness.
We propose to assess the human likeness of counterspeech based on the distinguishability between AI-generated and human-written replies to hate speech. To understand factors related to human likeness, we further analyze the differences between AI-generated and human-written counterspeech using linguistic factors, politeness, and specificity. Polite responses are essential for productive online communications Hermoyo et al. (2023). Specificity refers to targeted responses to hate speech that fit the conversation context, which is usually well captured by humans and vital for effective counterspeech Zheng et al. (2023).
This study addresses the following questions:
  • Which LLMs are better at mimicking human-written counterspeech?
  • Are there linguistic differences between AI-generated and human-written counterspeech?
  • Do AI-generated and human-written counterspeech differ in politeness and specificity?
We implement several state-of-the-art counterspeech generation strategies with LLMs, including prompting for one counterpeech reply(Prompt), prompt for multiple replies and select the best(Prompt and Select), fine-tune LLMs with existing counterspeech datasets(Fine-tune), and outcome-constrained LLMs.(Constrained). We also collect human-written counterspeech, including both examples from actual user-generated content (i.e., genuine content from Reddit) and crowd workers (i.e., counterspeech generated on demand).
Armed with AI-generated and human-written counterspeech, we build authorship attribution models to identify which AI models generate more human-like counterspeech. The higher the classification accuracy, the less human-like the AI-generated counterspeech is. We further deploy a human annotation task to validate and check whether humans can discern the differences between AI-generated and human-written counterspeech. Additionally, we analyze linguistic characteristics, politeness, and specificity of the counterspeech generated by various methods to understand the differences amongst counterspeech depending on the source (types of AI-generated and human-written).
Our study shows that AI-generated and human-written counterspeech can be easily distinguished by both classifiers and humans. Additionally, AI-generated counterspeech shows (a) significant differences in linguistic characteristics, and (b) is more polite and less specific.
Prior work Dataset Model Analysis
Domain Size Context Generation Model Training Methods
Gagiano and Tian (2023)
Translation;
News summarization;
Web text
Human: 9,000;
AI: 9,000
No T5, GPT-X Not specified Authorship classification
Zhou et al. (2023)
News writing;
Social media
Human: 12,408;
AI: 500
No GPT-3 Few-shot learning
Linguistic features;
Authorship classification
An et al. (2023) Scientific writing
Human: 400;
AI: 400
No ChatGPT Fine-tuning
Linguistic features;
Similarity comparison;
Authorship classification
Buz et al. (2024) Social media
Human: 411,189;
AI: 6,000
No
GPT-2,
GPT-Neo,
GPT-3.5-turbo
Zero-shot learning;
Fine-tuning
Linguistic features;
Human preference;
Authorship classification
Prajapati et al. (2024)
Scientific writing;
Social media; QA
Human: 30,482;
AI: 18,308
Yes ChatGPT-3 Zero-shot learning
Authorship classification;
Statistical imbalance;
Linguistic features;
Fact verification
Ours
HS/counterspeech;
Social media
Human: 29,181;
AI: 54,136
Yes Llama2
Prompt; Prompt and Select;
Fine-tune; Constrained
Human likeness;
Linguistic features;
Politeness; Specificity
Table 2: Comparison of differentiating AI-generated and human-written texts: An overview of previous research and our contributions. We are the first to explore several AI-generation strategies and target human likeness along with politeness and specificity.
2 Related Work
Counterspeech Generation Several studies employ crowd workers or experts to curate counterspeech Chung et al. (2019); Qian et al. (2019); Yu et al. (2022). These human-written datasets have been widely utilized for training counterspeech models in subsequent research Tekiroglu et al. (2022); Saha et al. (2022); Gupta et al. (2023). Hybrid approaches that combine human annotations and generative models have been proposed, resulting in counterspeech corpora such as CONAN Chung et al. (2019) and multiCONAN Fanton et al. (2021). Advanced generative models have been developed for more sophisticated generation of counterspeech, for example, the prompt and select method has been shown to generate more diverse and relevant counterspeech Zhu and Bhat (2021). Other approaches include constraining models for generating polite Saha et al. (2022), intent-based Gupta et al. (2023), or outcome-oriented counterspeech Hong et al. (2024). This study (a) implements several representative models to generate counterspeech and (b) conducts an extensive evaluation focusing on human likeness.
Counterspeech Evaluation Counterspeech evaluations have focused on relevance using BLEU, ROUGE  Chung et al. (2019); Qian et al. (2019); diversity with metrics like repetition rate  Zhu and Bhat (2021); Tekiroglu et al. (2022); and linguistic quality using GRUENV metrics Jiang et al. (2023a). These metrics offer insights limited to the lexical and semantic levels. Recently, researchers have developed evaluation metrics specifically tailored for counterspeech, such as politeness and effectiveness Saha et al. (2022); Hong et al. (2024). However, few studies have investigated the human likeness of AI-generated counterspeech. This study fills this void by evaluating several generation methods.
Comparing AI-Generated and Human-Written Texts Many studies have explored methods to differentiate AI-generated from human-written texts across domains and tasks. Examples include translation El-Sayed and Nasr (2023), news summarization Gagiano and Tian (2023), scientific writing Ma et al. (2023), and social media posts Buz et al. (2024). Recent studies utilize datasets ranging from a few hundred to over 40,000 samples. They employ models such as T5, GPT variants, and ChatGPT with zero-shot learning and fine-tuning. Table 2 presents a summary of these studies and the differences with the work presented here.
Most curated datasets consist of texts without additional context El-Sayed and Nasr (2023); Buz et al. (2024); An et al. (2023). Few include context and involve sequences or groups of related texts Ji et al. (2024); Prajapati et al. (2024). To distinguish between AI-generated and human-written texts, many researchers conduct linguistic analyses and authorship attribution, supplemented with human evaluations Zhou et al. (2023); Ma et al. (2023); Buz et al. (2024); Ji et al. (2024); Prajapati et al. (2024). Our study is among the few to work in a dialogue setting, where texts are dialogue turns countering a hate speech—AI-generated following four strategies or human-written from two sources.
3 Counterspeech Data Curation
Our experiments are grounded on large collections of hate speech posts and their counterspeech replies. The curation strategy includes both human-written counterspeech and AI-generated.
3.1 Human-written Counterspeech
Human-written counterspeech consists of replies to hateful content written by humans. We consider counterspeech written by both genuine Reddit users (i.e., users that post out of their own will) and crowd workers who are tasked with writing counterspeech. Genuine human-written counterspeech is collected from Reddit. First, we use keyword- and community-based sampling methods to retrieve hate speech posts from 42 subreddits (Appendix A) with a higher prevalence of hate via the Pushshift API. This step results in 27,491 hate speech posts and their replies. Second, we identify replies that are counterspeech with three BERT classifiers individually fine-tuned with three existing counterspeech corpora Qian et al. (2019); Chung et al. (2021); Yu et al. (2022). We consider a reply to be counterspeech if the three classifiers indicate so. This step results in 14,973 (hate speech, counterspeech) pairs from Reddit.
Counterspeech written by crow workers is collected from the Benchmark dataset Qian et al. (2019). This corpus includes hate speech posts paired with counterspeech replies written by crowd workers on demand. We identify 14,208 valid (hate speech, counterspeech) pairs in Benchmark.
3.2 AI-generated Counterspeech
We implement state-of-the-art strategies to generate counterspeech replies to hateful posts. The starting point is the 29,181 (hate speech, counterspeech) pairs from Section 3.1. Note that (a) some strategies require only hate speech posts while others also require the counterspeech reply and (b) all hate speech posts were written by real Reddit users.
We develop counterspeech generation models based on the following methods:
Prompt
LLMs are prompted to generate counterspeech given a hateful post Fraser et al. (2023); Hassan and Alikhani (2023); Saha et al. (2024).
Prompt and Select
LLMs are prompted to generate multiple counterspeech replies and classifiers are employed to select the most diverse and relevant one Zhu and Bhat (2021).
Fine-tune
LLMs are trained with (hate speech, counterspeech) pairs to learn to generate human-written counterspeech Chung et al. (2021); Fanton et al. (2021).
Constrained
We adopt reinforcement learning with LLMs for outcome-constrained generation Hong et al. (2024).
The last two strategies use the pairs described in Section 3.1 combined with three existing corpora: CONAN Chung et al. (2021), MultiCONAN Fanton et al. (2021), and Benchmark Qian et al. (2019). We present details (specific prompts, hyperparameters, etc.) in Appendix B. Ultimately, we obtain 54,136 AI-generated counterspeech replies (14,799, 19,436 14,215, and 5,686 respectively, as LLMs sometimes refuse to complete the task).
Prompt Prompt and Select Constrained Fine-tune
Agreement Rate 97% 97% 97% 94%
Cohen $\kappa$ 0.94 0.94 0.93 0.88
Table 3: Inter-annotator agreements differentiating human-written and AI-generated counterspeech. The task is straightforward for humans
4 Evaluation Methods
We conduct evaluations comparing human-written and AI-generated counterspeech accounting for four major categories: human likeness, linguistic differences, politeness and specificity.
Human Likeness
If it is easy to differentiate between human-written and AI-generated counterspeech, we can conclude that the latter is not human-like. To this end, we develop BERT classifiers with our curated dataset (Section 3). The dataset is divided into an 80/20 split for training and testing. Specifically, we create five binary classifiers (human-written or AI-generated) depending on what is included in the AI-generated counterspeech: all or only the counterspeech generated by one of the four strategies (see Appendix B for details).
In addition, we conduct a human validation for deeper insights. We randomly select 100 samples from each AI strategy (4×100)4$\times$ 100)4 × 100 ) and human-written counterspeech (2×100)2$\times$ 100)2 × 100 ), combine them, and ask human annotators to label whether the counterspeech has human-written or AI-generated. Two research assistants complete the annotation process. Table 3 presents the agreement rate and Cohen’s Kappa score. Both indicate high reliability, meaning that the task is straightforward for humans.
Linguistic Differences
We use SEANCE Crossley et al. (2017) to analyze the linguistic features of counterspeech and conduct statistical tests to reveal the differences between human-written and AI-generated counterspeech. For each type of AI-based method, we randomly sample an equal number of human-written counterspeech for comparison. We utilize the Wilcoxon rank-sum test to discern significant distinctions between the generation of AI and humans. Additionally, we apply the Bonferroni correction to identify the most significant linguistic features Weisstein (2004).
AI-generated Human-written
Prompt Prompt and Select Constrained Fine-tune User Crowd
Weighted Cohen $\kappa$ 0.80 0.94 0.96 0.93 0.80 0.88
Table 4: Inter-annotator agreements assessing politeness of AI-generated and human-written counterspeech.
Politeness
The level of politeness assesses the degree of respectfulness and courtesy. It provides an complementary perspective to understand the differences between AI-generated and human-written counterspeech. We build a politeness prediction model, a BERT model fine-tuned with the dataset by Saha et al. (2022). The model achieves an F1 score of 0.91. We use this model to predict the politeness level of the counterspeech replies on a scale of 0 to 7. A higher score indicates more politeness.
Additionally, we conduct a human validation to gain insights. We randomly select 100 samples from each AI strategy (4×100)4$\times$ 100)4 × 100 ) and human-written counterspeech (2×100)2$\times$ 100)2 × 100 ), combine them, and ask the same human annotators to label them on a scale from 0 to 7 (See Appendix C for the guidelines). Given the challenge of achieving exact agreement on a 7-point scale, the weighted Cohen’s Kappa is employed to calculate inter-annotator agreement (Table 4). The agreements ($\kappa\geq 0.8$) again indicate high reliability. We calculate the average of the human-annotated politeness scores and conduct the Kruskal-Wallis test to explore whether there are significant differences in the politeness of human-written and AI-generated counterspeech.
AI-generated Human-written
Prompt Prompt and Select Constrained Fine-tune User Crowd
Weighted Cohen $\kappa$ 0.87 0.86 0.80 0.95 0.93 0.81
Table 5: Inter-annotator agreements assessing the specificity of AI-generated and human-written counterspeech.
Specificity
We conduct a human assessment of the specificity of a counterspeech reply with respect to the corresponding hate speech post. This metric measures how well the counterspeech (a) aligns with contextual information and (b) targets the topic in the hate speech post Tekiroglu et al. (2022); Jones et al. (2024). We have designed a Likert scale (1–5) to evaluate specificity, where 5 indicates the highest specificity. In order to assess reliability, we randomly select 100 samples from each counterspeech source (700 in total) for annotators to assess. Inter-annotator agreement is again very high (Table 5, $\kappa\geq 0.80$).
5 Results
5.1 Human Likeness
Strategy AI Human Weighted Average
P R F1 P R F1 P R F1
Prompt 0.99 1.00 0.99 1.00 0.99 0.99 0.99 0.99 0.99
Prompt and Select 0.98 1.00 0.99 1.00 0.98 0.99 0.99 0.99 0.99
Fine-tune 0.80 0.95 0.87 0.94 0.76 0.84 0.87 0.86 0.86
Constrained 0.99 1.00 1.00 1.00 0.99 1.00 1.00 1.00 1.00
All 0.95 1.00 0.97 1.00 0.95 0.97 0.98 0.97 0.97
Table 6: Results of a BERT classifier differentiating (a) AI-generated counterspeech with each and all strategies and (b) human-written counterspeech. The high F1 scores indicate that AI-generated counterspeech is not human-like.
Classifier-Based Results Table 6 shows the results of the BERT-based classifier tuned to differentiate human-written and AI-generated counterspeech using the four strategies. We draw two main conclusions. First, the classifier performs well, achieving F1 scores above 0.97 for all strategies except Fine-tune. Second, the classifier struggles to distinguish counterspeech generated by Fine-tune and human-written counterspeech (weighted average F1: 0.86). These results indicate that the Fine-tune model more closely resembles human-written counterspeech compared to other generation models.
Corpus AI Human Weighted Average
P R F1 P R F1 P R F1
CONAN 0.97 0.98 0.97 0.98 0.96 0.97 0.97 0.97 0.97
Gab 0.86 0.99 0.92 0.99 0.84 0.91 0.93 0.92 0.92
Reddit 0.81 0.95 0.88 0.94 0.78 0.85 0.88 0.87 0.87
MultiCONAN 0.92 0.99 0.95 0.99 0.91 0.95 0.95 0.95 0.95
Effectiveness 0.96 0.98 0.97 0.98 0.96 0.97 0.97 0.97 0.97
Table 7: Results of a BERT classifier differentiating (a) AI-generated counterspeech with the Fine-tune strategy using different corpora and (b) human-written counterspeech. The corpora include CONAN Chung et al. (2019), MultiCONAN Fanton et al. (2021), Gab and Reddit from Benchmark Qian et al. (2019), and effective-oriented counterspeech Hong et al. (2024).
We further develop five classification models to investigate the differences when the Fine-tune strategy uses different datasets. Table 7 shows the results. The counterspeech generated by models fine-tuned with the Reddit benchmark dataset presents the greatest challenge. The weighted F1 score (0.87) is lower than others, however, it is still high. This leads to the observation that fine-tuned LLMs more closely capture the style of human-written counterspeech in the target platform—Reddit in this study. However, they are still limited in generating counterspeech that has a high human likeness. Models trained with expert-generated (i.e., CONAN, MultiCONAN), or outcome-oriented (i.e., effectiveness) counterspeech are the easiest to differentiate (F1: 0.97, 0.95, and 0.97, respectively.
Strategies AI Human Weight Average
P R F1 P R F1 P R F1
Prompt 0.91 1.00 0.95 1.00 0.93 0.97 0.96 0.96 0.96
Prompt and Select 1.00 0.92 0.96 0.93 1.00 0.96 0.96 0.96 0.96
Constrained 0.90 0.98 0.94 0.98 0.91 0.94 0.94 0.94 0.94
Fine-tune 0.49 0.64 0.55 0.63 0.48 0.55 0.57 0.55 0.55
Table 8: Results obtained by humans differentiating (a) AI-generated counterspeech with each strategy and (b) human-written counterspeech. Humans are much less reliable than the BERT classifier (Table 6 ) identifying counterspeech obtained with the Fine-tune strategy, but otherwise are proficient.
Human Assessment Table 8 presents the results when humans differentiate AI-generated and human-written counterspeech. Counterspeech generated by Prompt, Prompt and Select, and Constrained can be easily identified by human annotators. The counterspeech by Fine-tune, however, is more challenging to distinguish, a consistency mirrored in the computing-based evaluation.
Human performance is lower than model performance across all AI-generation methods. We conduct an error analysis and find the following. First, Some human-written counterspeech is template-based, for example, “Use of such language or words is not acceptable” or “Using that language doesn’t help you make your point.” Though counterspeech examples are human-written, their formulaic nature resembles robotic outputs, making it challenging for human annotators to distinguish them from AI-generated. Second, Fine-tune models can better mimic human-like responses, generating responses that closely resemble authentic replies on Reddit. This kind of generation mimics human emotion and language style, complicating the task of differentiating it from genuine human-written counterspeech.
Category Prompt Prompt and Select Constrained Fine-tune All
Textual factors
[…]
Emotional factors
[…]
Social factors
[…]
Table 9: Linguistic analysis comparing counterspeech generated by AI and humans across different AI-based generation methods. The up arrow indicates higher values in AI-generated counterspeech. The number of arrows indicates the p-value of the Wilcoxon rank-sum test (one: $p<0.05$, two: $p<0.01$, and three: $p<0.001$). All tests have passed Bonferroni correction.
5.2 Linguistic Differences
We summarize the findings into three factors: textual, emotional, and social in Table 9. Various linguistic differences exist between human-written and AI-generated counterspeech. The trend is consistent across most groups except Fine-tune. Counterspeech generated by Fine-tune tends to be more human-like, exhibiting distinct linguistic patterns compared to other AI-generated groups.
Textual factors refer to the language style, structure, and function. Human-written counterspeech tends to contain more words of action, format (except Constrained), frequency, and overstated (except Fine-tune). AI-generated counterspeech includes more 1st person pronouns (except Prompt), self-expression words, and anticipation words (except Fine-tune), which is consistent with the insight of  Muñoz-Ortiz et al. (2023). The use of certainty words is significantly higher in AI-generated counterspeech  Zhou et al. (2023), and this trend is consistent across all groups.
Regarding emotional factors, human-written counterspeech expresses stronger feelings, such as negative, hatred, and excitement. AI-generated counterspeech significantly conveys more positive emotion (except Fine-tune).
Social factors are associated with social dynamics, norms, values, and structures, referring to the social context of the content. Most social factors are prevalent in the human-written counterspeech, including religious, respect, wealth, and power words. AI-generated counterspeech tends to contain more economic words (except Fine-tune).
Figure 1: Politeness scores distribution of human-written and AI-generated counterspeech. Higher values indicate more politeness. AI-generated counterspeech is more polite.
5.3 Politeness
Classifier-Based Results We conduct the Wilcoxon rank sum test between human-written and AI-generated counterspeech. The results show that the politeness of human-written counterspeech is significantly lower than AI-generated ($p<0.001$, Mean: 2.14 vs 2.37). Figure 1 presents the politeness distribution of all the human-written and AI-generated counterspeech.
Both types of generation have high density in lower politeness scores, with approximately 25% at 0 and 35% at 1. However, human-written counterspeech tends to concentrate on lower politeness levels than AI-generated counterspeech. AI-generated counterspeech has a higher proportion in higher politeness scores, especially at a score of 6 (human-written: 12%, AI-generated: 20%).
Figure 2(a) shows the human annotation results of politeness. The results align with the findings of the evaluation using the classifier: AI-generated counterspeech demonstrates notably higher levels of politeness than human-written counterspeech. Responses by social media users are less polite than crowdsourced and AI-generated counterspeech.
In AI-generated methods, the Constrained method generates more polite responses, with significantly higher polite scores according to the Kruskal-Wallis test ($p<0.001$). The counterspeech generated by Fine-tune exhibits a notable spread towards both high and low ends, suggesting that counterspeech can range from very polite to impolite. Counterspeech in groups of Prompt and Select is less polite than that from Fine-tune and Constrained, but is still notably more polite than the user and crowdsourcing generation.
5.4 Specificity
Figure 2: Distribution of politeness (left) and specificity (right) scores assigned by humans. We plot the distributions per strategy to generate counterspeech (four left-most plots) and human source (two right-most plots)).
The counterspeech generated by Prompt and Prompt and Select models exhibits a similar distribution with a moderate specificity level of 3 (Figure 2(b)). The Constrained model-generated counterspeech has a median specificity score of 1.5, indicating that it lacks specificity and provides a more general reply. Comparatively, the counterspeech by Fine-tune models shows high variances in the specificity scores, but most scores are between 3 and 4, indicating counterspeech more targeted to the hate speech post at hand.
Crowdsourced counterspeech shows lower specificity(median score: 2.5). We find that a lot of crowdsourced counterspeech follows templates, uses vague wording, and does not directly address the specific hate speech post at hand. The user-generated counterspeech exhibits a broad distribution ranging from 1 to 5, with a higher frequency in scores of 4 and 5, suggesting some counterspeech has good specificity. Annotators also find that user-generated counterspeech is more contextually relevant and better targets the issues in a hateful post.
6 Discussion
According to evaluations in this study, AI-generated counterspeech can be easily differentiated by classification models and humans. Comparatively, Fine-tune models outperform other automatic strategies in generating more human-like responses. These responses more closely resemble human-written counterspeech in linguistic features, exhibit a wide range of politeness levels, and have higher specificity scores.
Counterspeech generated by Prompt and Prompt and Select tend to be long and lose efficacy. Some generic counterspeech (e.g., “Finally, I would like to encourage you to engage in constructive conversation […]”) is commonplace. While the response is well-intentioned, it does not address the specific content of the hateful post. Additionally, counterspeech by Prompt and Prompt and Select also shows template expressions such as “It’s understandable that […]” and “By working together to address these issues, we can build a society […]”, which are generic.
Counterspeech by Constrained is significantly more polite, with expressions like: “ I apologize […] I would encourage you to […]”. But there is a lot of repetitive content in the counterspeech, including “I encourage you to strive for inclusivity, empathy, and respect for all people.” Such responses seem template-based and can be easily distinguished from human-written counterspeech.
In AI-generated counterspeech, we observe a common limitation: some counterspeech focuses on criticizing specific words and often includes misunderstandings. For example, terms such as “retard” are not always aimed at individuals with mental disabilities, yet counterspeech frequently responds with statements like “Using a word that describes someone with a mental disability does not promote understanding.” This type of counterspeech can be awkward or counterproductive.
Crowdsourced counterspeech often follows a template, exhibiting a limited range of language styles. Many responses are not contextual and can be applied to various scenarios (e.g., “Use of such language or words is not acceptable.”) It indicates that when humans are paid to generate counterspeech, they may repeat template-based replies that come across as repetitive or even robotic.
User-written counterspeech demonstrates higher specificity, which can better capture the nuances in hate speech, offering more targeted and related responses, resonating the findings by  Go and Sundar (2019a). Since humans are better at interpreting human communication, they can more precisely grasp the intent behind hate speech Chen et al. (2018). However, user-written counterspeech may be less polite. It is common for human response to convey natural emotions including anger, toxicity, and impoliteness when expressing opinions.
7 Conclusion
We propose to evaluate the human likeness of AI-generated counterspeech. We implement state-of-the-art counterspeech generation models following four strategies (Prompt, Prompt and Select, Fine-tune, and Constrained), and use a Reddit hate speech / counterspeech dataset Qian et al. (2019) for counterspeech generation and evaluation. The human-written counterspeech comprises responses by crowd workers and social media users.
We perform evaluations in authorship identification, linguistic features, politeness, and specificity. We find that counterspeech generated by current state-of-the-art models is distinguishable by both algorithms and humans. There are significant differences between AI-generated and human-written counterspeech in linguistic features, politeness, and specificity. AI-generated counterspeech is more polite and less focused, which are potential factors that make them differentiable from human-generated counterspeech. Fine-tuning LLMs with pertinent datasets makes the counterspeech generation more human-like. A future research direction might be the development of LLMs for generating more human-like replies.
Limitations
Our study has some limitations: (1) Not including all counterspeech generation methods. We focus on four popular generation methods (Prompt, Prompt and Select, Constrained, and Fine-tune) which are predominant in recent research. However, we do not explore other counterspeech generation strategies due to time and cost constraints. Nonetheless, our evaluation methods can be applied to assess other generation methods. (2) Limited evaluation scope. While our evaluation attempts to capture the nuances of human-written and AI-generated counterspeech, there are aspects such as tone and cultural sensitivity that are not evaluated, presenting avenues for future research. (3) Subjectivity in human annotation. The process of annotating human likeness and politeness involves a degree of subjectivity that can lead to variability in results. We point out, however, that our human assessments follow standards for high reliability (double annotation and high agreements). (4) Lack of cross-domain generalization. We focus on Reddit data, comparing AI-generated counterspeech with crowdsourcing counterspeech and user-generated counterspeech from Reddit. The results may vary depending on the nature of the original content and the platform on which it is applied. (5) Immediate research is needed. As AI models are constantly improving, the performance of AI-generated results might change over time. In this study, we take the relatively new LLM Llama2 models for evaluation, however. More work needs to be done to identify whether our findings hold true for other models and newer versions.
Ethics Statement
We fully consider the potential risks and benefits of our study. First, we collect data from Reddit, a public forum that makes data available to third parties. We have masked users’ names and identities before analysis. Second, our study employs human annotators to evaluate counterspeech, their names and identities are encrypted to avoid the identification of annotators. Third, we provide various AI-based methods to generate counterspeech that embody human-likeness. There is a concern that some generated responses are indistinguishable from those created by humans, raising ethical considerations. We acknowledge the potential risks these methods may cause. However, the benefits will outweigh such risks, for example, using these generations in adversarial learning to develop more robust models for identifying AI-generated content.
Acknowledgements
We acknowledge the use of code-writing assistance GitHub Copilot. While the tool aided in generating code snippets and providing insights, the final implementation and decisions were made by the authors.
Appendix A Subreddit List
We collect Reddit data from the following 42 subreddits (Table 10).
Subreddits
r/antiwork, r/changemyview, r/NoFap, r/Seduction,
r/PurplePillDebate, r/ShitPoliticsSays, r/PurplePillDebate,
r/bindingofisaac, r/FemaleDatingStrategy,r/SubredditDrama
r/KotakuInAction, r/DotA2, r/technology, r/modernwarfare,
r/playrust, r/oblivion
r/bakchodi, r/Feminism, r/PussyPass, r/MensRights,
r/Sino, r/BlackPeopleTwitter, r/india, r/PussyPassDenied,
r/TwoXChromosomes, r/GenZedong, r/antheism
r/4Chan, r/justneckbeardthings, r/HermanCainAward,
r/MetaCanada, r/DankMemes, r/ShitRedditSays
r/conspiracy,r/worldnews, r/Drama, r/TumblrInAction,
r/lmGoingToHellForThis, r/TrueReddit.
Table 10: Subreddit List.
Appendix B Generation Details
We implement several prominent generation methods in our study, categorized into four types: Prompt, Select, Constrained, and Fine-tune. We share the details of our experimental models.
Model Parameters Due to limited computing resources, we use the Llama-2-7b-chat model for all generation experiments. This is to minimize the potential impact of using different LLMs on the generation results, thereby allowing the comparison to better reflect the differences in training methods. Specifically, the top k is set to be 8, the temperature is 0.7, and the maximum length of reply is 512.
Expertment details Prompt: We request the Llama model to generate counterspeech based on our designed prompt in the following. This prompt is used in all generation methods.
System: "Generate a response in Reddit Style." User: "Here is the Reddit comment: <Hate Comment>. Please write a counterspeech to the Reddit hate comment."
The LLMs may avoid answering inappropriate questions and produce null results. We exclude null results in the evaluation.
Select: We request the model to generate 10 different responses for each HS, then implement the pruning and selection method by  Zhu and Bhat (2021) to generate counterspeech.
Finetune: We fine-tune the Llama-2-7b-chat model separately with the CONAN Chung et al. (2019), MultiCONAN Fanton et al. (2021), Gab, and Reddit data from Benchmark Qian et al. (2019). The Finetune models can generate short responses. We remove responses with fewer than four words for further evaluation.
Constrarined: We incorporate reinforcement learning to contain the counterspeech generation following Hong et al. (2024). We design reward models to guide the reinforcement learning process: the constructive conversation outcome classifier Yu et al. (2024). We first trained a finetuned model with the Reddit data from Benchmark to control the initial generation process, ensuring the model can generate responses to HS. Then the model is further trained with the reinforcement learning process, where responses that lead to low conversation incivility are awarded higher. The model after the reinforcement learning process is then used for counterspeech generation.
Computing Resources The computational resources applied in this research include a high-performance server equipped with an Intel Xeon Gold 6226R processor, 128 GB memory, and 3 Nvidia RTX 8000 GPUs.
Appendix C Annoatation Details
We employ two PhD students with expertise in counterspeech study for the annotation tasks. Annotators were compensated on average with \$15 per hour. Annotators are females, one is a US citizen and one is an international student.
To maintain high-quality annotation in our study, we provide annotation training before human assessment. We have randomly collected 20 examples from the dataset for practice for each task. We require the annotators to review the HS first and then assess the corresponding counterspeech. After they annotate the data, we will check the labeled score and have a discussion with them to revise their understanding. After several rounds, their annotations show a higher consistency with expert annotation, indicating the great quality of the work.
We also provide detailed annotation rubrics and examples to guide the human evaluation in Table 11. We have also designed the annotation rubric and examples for specificity in Table 12. Examples here may not represent the rubric very accurately, since it should be interpreted combined with HS. After two annotators complete their task, an expert reviews the results and makes final revisions. The final labels for human likeness are determined by the expert, while the final politeness and specificity scores are calculated as the average of the two annotators’ labels.
Score Rubric Example
0-Very Impolite Counterspeech is percieved as rude or unfriendly The study was b**.
1-Impolite Counterspeech that might not be overtly rude but lack warmth. Use of the r-word is unacceptable
2-Somewhat Impolite
Counterspeech is neutral but could include language that might
be perceived as slightly unfriendly.
I don’t think you’re understanding what I’m saying.
3-Neutral
Counterspeech that is politely formulated without additional
warmth or friendliness but not impolite.
He just doesn’t have the mental capacity to understand
the consequences of his actions
4-Somewhat Polite
Counterspeech that are polite and show an attempt at
engagement without significant warmth.
Please be careful with the words you choose!
5-Polite
Counterspeech that clearly make an effort to be
considerate and courteous.
Instead, I suggest that you strive to promote inclusivity,
empathy, and understanding towards all individuals
6-Very Polite
Counterspeech includes expressions that enhance the
interaction positively.
It’s not okay to use derogatory language or slurs
to describe any group of people, Everyone deserves to
be treated with respect and dignity, regardless of their
beliefs or actions.
7-Extremely Polite Counterspeech is very thoughtful or considerate.
Please refrain from using language that is
inappropriate or offensive, as it is not productive
or respectful. Thank you. Is there something else
I can help you with?
Table 11: Politeness Rubric and Examples.
Score Rubric Example
1-Very Low Specificity
Counterspeech is generic and does not address the specific content
or context of HS.
It’s offensive and hurtful to people with actual mental disabilities.
2-Low Specificity Counterspeech addresses the general theme of HS but lacks details. Please refrain from using hateful ableist language in your posts
3-Moderate Specificity
Counterspeech somewhat addresses specific aspects of HS,
but may still include some general statements
I think they mean that white people have an advantage in society.
4-High Specificity
Counterspeech directly engages with specific points HS,
providing targeted rebuttals.
I’m not being retarded. I’m just not understanding why it’s bad.
5-Very High Specificity Counterspeech is highly focused and precisely counters HS.
The government is not in the business of entitling people to free passes.
The government is in the business of protecting people’s rights.
Table 12: Specificity Rubric and Examples.

### v2

@@
 Counterspeech is a targeted response to counteract and challenge abusive or hateful content.
-It can effectively curb the spread of hatred and foster constructive online communication.
+It effectively curbs the spread of hatred and fosters constructive online communication.
 Previous studies have proposed different strategies for automatically generated counterspeech.
-Evaluations, however, focus on the relevance, surface form, and other shallow linguistic characteristics.
-In this paper, we investigate the human likeness of AI-generated counterspeech, a critical factor influencing effectiveness.
+Evaluations, however, focus on relevance, surface form, and other shallow linguistic characteristics.
+This paper investigates the human likeness of AI-generated counterspeech, a critical factor influencing effectiveness.
 We implement and evaluate several LLM-based generation strategies, and discover that AI-generated and human-written counterspeech can be easily distinguished by both simple classifiers and humans.
-Further, we reveal differences in linguistic characteristics, politeness and specificity.
+Further, we reveal differences in linguistic characteristics, politeness, and specificity.
+The dataset used in this study is publicly available for further research  [^1] https://github.com/oliveeeee25/counterspeech_eval_humanlike.
 Assessing the Human Likeness of AI-Generated Counterspeech
-Xiaoying Song1, Sujana Mamidisetty1, Eduardo Blanco2, Lingzi Hong1 1College of Information, University of North Texas 2Department of Computer Science, University of Arizona
-Triggering warning: Read with caution. Examples in this paper may present toxic languages.
+Xiaoying Song1 Sujana Mamidisetty2 Eduardo Blanco3 Lingzi Hong1 1College of Information, University of North Texas 2TAMS, University of North Texas 3Department of Computer Science, University of Arizona {xiaoyingsong, sujanamamidisetty}@my.unt.edu eduardoblanco@arizona.edu lingzi.hong@unt.edu
+Trigger warning: Read with caution. Examples in this paper may present toxic languages.
 1 Introduction
@@
 I suggest that we focus on promoting empathy, understanding, and constructive dialogue.
-Table 1: A hate speech post from Reddit and AI-generated counterspeech reply.
-In this paper, we compare AI-generated and human-written counterspeech replies and conclude that are fundamentally different.
-Research has been devoted to curating counterspeech Qian et al.
+Table 1: A hate speech post from Reddit and counterspeech generated by prompt with Llama2. We compare AI-generated and human-written counterspeech and conclude they are fundamentally different.
+Recent studies have explored methods for generating counterspeech Qian et al.
 (2019); Chung et al.
-(2021); Yu et al.
 (2021).
-Human-written counterspeech can positively affect bystanders and encourage more civil online interaction  Zhao et al.
+Human-written counterspeech is often contextually appropriate and can positively affect bystanders and encourage civil online interactions  Zhao et al.
 (2024).
-However, human involvement is often time-consuming and costly.
-Generative AI models have also been developed to automatically generate counterspeech Zhu and Bhat (2021); Tekiroglu et al.
-(2022); Saha et al.
-(2022); Hassan and Alikhani (2023); Hong et al.
+However, human involvement is time-consuming and costly.
+Models have been developed to automatically generate counterspeech Zhu and Bhat (2021); Tekiroglu et al.
+(2022); Hong et al.
 (2024).
-Automatic generation has the potential to mimic human-written counterspeech and assist in combating hate Bail (2024).
-However, generative AI models may struggle to accurately understand semantic nuances, leading to misunderstandings and even backfiring Luger and Sellen (2016); Hadi (2019); Toader et al.
+The provision has the potential to mimic human-written counterspeech and assist in combating hate Bail (2024).
+However, counterspeech generation models may struggle to accurately understand semantic nuances, leading to misunderstandings and even backfiring Luger and Sellen (2016); Hadi (2019); Toader et al.
 (2019).
-While these systems attempt to convey empathy like humans, sometimes they do not appropriately match the user’s emotional state or the conversation context  Han et al.
+Sometimes they do not appropriately match the user’s emotional state or the conversational context  Han et al.
 (2022); Chung et al.
@@
 Table 1 shows an example of counterspeech generated by prompt with Llama2.
-While it aims to mimic human responses, the generated counterspeech may lack the depth of empathy and cultural sensitivity crucial for effective communication Benesch et al.
-(2016); Hangartner et al.
+The counterspeech is perceived to be robotic, as it ignores the hatred and anger directed at the victim in the language (“racist feminist c*** …”) and instead focuses on criticizing the imagined scenario (“not put it out if she were on fire”).
+The final suggestion is vague and does not specifically address the harmful stereotypes in the original statement.
+It may lack nuanced understanding and the depth of empathy for effective communication Hangartner et al.
 (2021).
@@
 (2024).
-The tone and form of counterspeech can significantly affect the effectiveness in influencing behaviors of users Buerger (2021).
-Generated counterspeech that embodies human characteristics including empathy, humor, and respect can foster emotional attachments among users Glikson and Woolley (2020), potentially making counterspeech more acceptable and reliable Go and Sundar (2019a); Jiang et al.
+The tone and form of counterspeech significantly affect the effectiveness in influencing behaviors of users Buerger (2021).
+Generated counterspeech that embodies human characteristics including empathy, humor, and respect can foster emotional attachments Glikson and Woolley (2020), potentially making counterspeech more acceptable and reliable Jiang et al.
 (2023b).
-Human-like responses are more likely to express understanding and social connection Go and Sundar (2019a), contribute to fostering genuine interactions, and build trust Fenwick and Molnar (2022); Cheng and Pan (2021).
+Human-like responses are more likely to express understanding and social connection Go and Sundar (2019a), contribute to fostering genuine interactions, and build trust Fenwick and Molnar (2022).
 On the other hand, there are ethical concerns if AI-generated and human-written counterspeech are indistinguishable Hua et al.
@@
 AI models may be used for manipulative purposes such as swaying opinions not aligned with specific agendas Karnouskos (2020).
-AI models may also generate biased responses that could unintentionally reinforce stereotypes or cause harm Lucy and Bamman (2021); Kenthapadi et al.
+These models may also generate biased responses that could unintentionally reinforce stereotypes or cause harm Lucy and Bamman (2021); Kenthapadi et al.
 (2023).
@@
 Previous studies have focused on evaluating the relevance, quality, toxicity, persuasiveness, and effectiveness of counterspeech Saha et al.
-(2022); Jiang et al.
-(2023a); Hong et al.
+(2022); Hong et al.
 (2024).
-They have also studied linguistic aspects using metrics such as BLEU, METEOR, BERTScore, GRUENV scores, type-token and distinct-n Chung et al.
-(2019); Qian et al.
-(2019); Chung et al.
-(2021); Zhu and Bhat (2021); Tekiroglu et al.
-(2022); Jiang et al.
-(2023a).
+Some studied linguistic aspects using metrics such as BLEU, METEOR, BERTScore, GRUENV scores, type-token and distinct-n Zhu and Bhat (2021); Tekiroglu et al.
+(2022).
 However, few have conducted thorough evaluations regarding human likeness.
@@
   • Do AI-generated and human-written counterspeech differ in politeness and specificity?
-We implement several state-of-the-art counterspeech generation strategies with LLMs, including prompting for one counterpeech reply(Prompt), prompt for multiple replies and select the best(Prompt and Select), fine-tune LLMs with existing counterspeech datasets(Fine-tune), and outcome-constrained LLMs.(Constrained).
+We implement several state-of-the-art counterspeech generation strategies with LLMs, including prompting for one counterspeech (Prompt), prompt for multiple replies and select the best(Prompt and Select), fine-tune LLMs with existing counterspeech datasets(Fine-tune), and outcome-constrained LLMs.(Constrained).
 We also collect human-written counterspeech, including both examples from actual user-generated content (i.e., genuine content from Reddit) and crowd workers (i.e., counterspeech generated on demand).
-Armed with AI-generated and human-written counterspeech, we build authorship attribution models to identify which AI models generate more human-like counterspeech.
+With AI-generated and human-written counterspeech, we build authorship attribution models to identify which AI models generate more human-like counterspeech.
 The higher the classification accuracy, the less human-like the AI-generated counterspeech is.
@@
 Additionally, we analyze linguistic characteristics, politeness, and specificity of the counterspeech generated by various methods to understand the differences amongst counterspeech depending on the source (types of AI-generated and human-written).
-Our study shows that AI-generated and human-written counterspeech can be easily distinguished by both classifiers and humans.
-Additionally, AI-generated counterspeech shows (a) significant differences in linguistic characteristics, and (b) is more polite and less specific.
+Our study shows both classifiers and humans can easily distinguish AI-generated and human-written counterspeech.
+AI-generated counterspeech shows (a) significant differences in linguistic characteristics, and (b) is more polite and less specific.
 Prior work Dataset Model Analysis
-Domain Size Context Generation Model Training Methods
+Domain Size Context? Generation Model Training Methods
 Gagiano and Tian (2023)
@@
 Ours
-HS/counterspeech;
+Hate speech/counterspeech;
 Social media
@@
 (2019); Qian et al.
-(2019); Yu et al.
 (2019).
-These human-written datasets have been widely utilized for training counterspeech models in subsequent research Tekiroglu et al.
+These human-written datasets have been widely utilized for training counterspeech models Tekiroglu et al.
 (2022); Saha et al.
@@
 (2021).
-Advanced generative models have been developed for more sophisticated generation of counterspeech, for example, the prompt and select method has been shown to generate more diverse and relevant counterspeech Zhu and Bhat (2021).
+Advanced generative models have been developed for a more sophisticated generation of counterspeech, for example, the prompt and select method has been shown to generate more diverse and relevant counterspeech Zhu and Bhat (2021).
 Other approaches include constraining models for generating polite Saha et al.
@@
 (2024).
-This study (a) implements several representative models to generate counterspeech and (b) conducts an extensive evaluation focusing on human likeness.
+This study implements several representative models to generate counterspeech and conducts an extensive evaluation focusing on human likeness.
 Counterspeech Evaluation Counterspeech evaluations have focused on relevance using BLEU, ROUGE  Chung et al.
@@
 (2024).
-To distinguish between AI-generated and human-written texts, many researchers conduct linguistic analyses and authorship attribution, supplemented with human evaluations Zhou et al.
+To distinguish between AI-generated and human-written texts, many researchers conduct linguistic analysis and authorship attribution, supplemented with human evaluations Zhou et al.
 (2023); Ma et al.
@@
 Our study is among the few to work in a dialogue setting, where texts are dialogue turns countering a hate speech—AI-generated following four strategies or human-written from two sources.
-3 Counterspeech Data Curation
+3 Counterspeech Data
 Our experiments are grounded on large collections of hate speech posts and their counterspeech replies. The curation strategy includes both human-written counterspeech and AI-generated.
@@
 Genuine human-written counterspeech is collected from Reddit.
-First, we use keyword- and community-based sampling methods to retrieve hate speech posts from 42 subreddits (Appendix A) with a higher prevalence of hate via the Pushshift API.
+First, we use keyword- and community-based sampling methods to retrieve hate speech posts from 42 subreddits (Appendix A) with a higher prevalence of hate via the Pushshift API  [^2] https://pushshift.io/api-parameters/.
 This step results in 27,491 hate speech posts and their replies.
@@
 (2022).
-We consider a reply to be counterspeech if the three classifiers indicate so.
-This step results in 14,973 (hate speech, counterspeech) pairs from Reddit.
-Counterspeech written by crow workers is collected from the Benchmark dataset Qian et al.
+We consider a reply to be counterspeech if the three classifiers indicate so, finally resulting in 14,973 (hate speech, counterspeech) pairs from Reddit.
+Counterspeech written by crowd workers is collected from the Benchmark dataset Qian et al.
 (2019).
@@
 4 Evaluation Methods
-We conduct evaluations comparing human-written and AI-generated counterspeech accounting for four major categories: human likeness, linguistic differences, politeness and specificity.
+We conduct evaluations comparing human-written and AI-generated counterspeech accounting for four major categories: human likeness, linguistic differences, politeness, and specificity.
 Human Likeness
 If it is easy to differentiate between human-written and AI-generated counterspeech, we can conclude that the latter is not human-like.
-To this end, we develop BERT classifiers with our curated dataset (Section 3).
+We develop BERT classifiers with our curated dataset (Section 3).
 The dataset is divided into an 80/20 split for training and testing.
@@
 In addition, we conduct a human validation for deeper insights.
-We randomly select 100 samples from each AI strategy (4×100)4$\times$ 100)4 × 100 ) and human-written counterspeech (2×100)2$\times$ 100)2 × 100 ), combine them, and ask human annotators to label whether the counterspeech has human-written or AI-generated.
+We randomly select 100 samples from each AI strategy (4×100)4$\times$ 100)4 × 100 ) and human-written counterspeech (2×100)2$\times$ 100)2 × 100 ), combine them, and ask human annotators to label whether the counterspeech is human-written or AI-generated.
 Two research assistants complete the annotation process.
@@
 The level of politeness assesses the degree of respectfulness and courtesy.
-It provides an complementary perspective to understand the differences between AI-generated and human-written counterspeech.
+It provides a complementary perspective to understand the differences between AI-generated and human-written counterspeech.
 We build a politeness prediction model, a BERT model fine-tuned with the dataset by Saha et al.
@@
 The Constrained model-generated counterspeech has a median specificity score of 1.5, indicating that it lacks specificity and provides a more general reply.
-Comparatively, the counterspeech by Fine-tune models shows high variances in the specificity scores, but most scores are between 3 and 4, indicating counterspeech more targeted to the hate speech post at hand.
+Comparatively, the counterspeech by Fine-tune models shows high variances in the specificity scores, but most scores are between 3 and 4, indicating counterspeech is more targeted to the hate speech post at hand.
 Crowdsourced counterspeech shows lower specificity(median score: 2.5).
@@
 Second, our study employs human annotators to evaluate counterspeech, their names and identities are encrypted to avoid the identification of annotators.
-Third, we provide various AI-based methods to generate counterspeech that embody human-likeness.
+Third, we provide various AI-based methods to generate counterspeech that embody human likeness.
 There is a concern that some generated responses are indistinguishable from those created by humans, raising ethical considerations.
@@
 Acknowledgements
-We acknowledge the use of code-writing assistance GitHub Copilot.
-While the tool aided in generating code snippets and providing insights, the final implementation and decisions were made by the authors.
+This work was supported by the Institute of Museum and Library Services (IMLS) National Leadership Grants under LG256661-OLS-24 and LG-256666-OLS-24.
 Appendix A Subreddit List
@@
 Score Rubric Example
-0-Very Impolite Counterspeech is percieved as rude or unfriendly The study was b**.
+0-Very Impolite Counterspeech is perceived as rude or unfriendly The study was b**.
 1-Impolite Counterspeech that might not be overtly rude but lack warmth. Use of the r-word is unacceptable